# The art of modeling nuclear reactions with weakly bound nuclei: status and perspectives

### v1

@@ 1 Introduction
Nuclear reactions are key tools to extract information about the structure of atomic nuclei, and the dynamical phenomena arising from the nucleus-nucleus forces. When one of the colliding partners is weakly bound, new phenomena and mechanisms arise, requiring in some cases specific reaction frameworks tailored to the peculiarities of these systems and their interactions.
Weakly-bound nuclei appear in the proximity to the neutron and proton driplines, a region where new exotic structures and phenomena are found. Promiment examples are halo nuclei, weakly bound nuclei composed of a compact core and one or two loosely bound nucleons with an unusually large matter radius, or Borromean systems, three-body systems with no bound binary subsystems, such as ${}^{9}\text{Be}~{}(\alpha+\alpha+n)$ or ${}^{6}\text{He}~{}(\alpha+n+n)$.
Although the field has experienced a great impulse in the last decades, the physics of nuclear scattering with weakly-bound projectiles is not new and many of the problems that are being addressed were already recognized much earlier as, for example, in the context of low-energy deuteron scattering. The deuteron, while being a stable nucleus, displays many halo-like features, such as weak binding ($E_{b}=2.22$ MeV), large spatial extension (the proton-neutron separation is about 3.8 fm) and no bound excited states. Let us consider the scattering of low-energy deuterons (a few MeV) by a heavy target nucleus, like 208Pb. Initially, when the deuteron is far apart from the target nucleus, it is in an internal state given by the ground-state of the proton-neutron Hamiltonian. As the deuteron approaches the target, it will feel its Coulomb repulsion. If the deuteron were a point-like particle, the effect of this interaction would be to distort the trajectory of the deuteron, without altering its internal structure. However, this is not the case. The Coulomb interaction acts on the proton, whose distance from the center-of-mass of the deuteron is $\sim$2 fm. Consequently, in addition to the monopole Coulomb potential ($\propto 1/R$) the deuteron will feel higher-order Coulomb multipoles arising from the expansion, valid for $r\ll R$, given by Cle62b
@@
where $\vec{R}$ is the coordinate from the target to the deuteron c.m.  $\vec{r}$ the proton-neutron relative coordinate and $\theta$ is the angle between them (see Fig. 1). The main deviation from the point Coulomb interaction is caused by the dipole term which depends on the orientation of the proton-neutron relative coordinate with respect to the deuteron-target coordinate. Specifically, when the proton is closer to the target nucleus ($\cos(\theta)<0$), the dipole potential will add a positive contribution, whereas when the proton is farther from the c.m. of the deuteron the contribution will be negative. Classically, we may think of this problem as an electric dipole moving in a slowly-varying electric field, in which the tidal force exerted on the dipole will favour the configuration of Fig. 1. Quantum-mechanically, the perturbation induced by the dipole force will couple the deuteron ground state (positive parity) with negative-parity states. Since the ground-state is the only bound state, these states necessarily appear in the continuum. Therefore, the polarization effect will modify the state of the system, producing a new one in which the ground state is mixed with continuum states of negative parity. This has a twofold effect on the outcome of the scattering process. First, it will produce deviations of the elastic scattering with respect to the Rutherford formula. Second, the coupling with the positive-energy states will give rise to some dissociation probability, that is, the breakup of the deuteron into a proton and a neutron.
The effect of dipole polarizability can be included by means of a polarization potential that, added to the monopole Coulomb potential, describes the motion of the center of mass of the nucleus. In the adiabatic limit, it is assumed that the excitation energies are high enough so the characteristic time for a transition to a state is small compared to the characteristic time for the collision. Applying second-order perturbation theory, one gets the following expression for this adiabatic polarization potential Ald75 :
$V_{\mathrm{ad}}(R)=-\sum_{n\neq 0}\frac{|\langle n|V_{\mathrm{dip}}|0\rangle|^{2}}{E_{n}-E_{0}}=-\frac{1}{2}\alpha\frac{(Z_{T}e)^{2}}{R^{4}},$ (2)
To conclude with this example, it is to be noted that the adiabiatic polarization potential given by Eq. (2) is purely real and, as such, does not describe the effect of the deuteron breakup that would remove flux from the elastic channel. This point will be revisited in Sec. 7.3, where another version of the polarization potential, which accounts for the deuteron breakup, will be presented.
The polarization potential adds an attractive contribution to the Coulomb point-particle potential that, when used in the Schrödinder equation, will produce a small, but measurable deviation of the elastic cross section with respect to the well-known Rutherford formula. An accurate measurement of sub-Coulomb elastic scattering data can therefore be used to extract the dipole polarizability of the deuteron (and other polarizable systems). The idea has in fact been successfully applied to the case of deuteron scattering Rod82 and constitutes a neat and beautiful example in which an important structure quantity can be inferred from reaction observables.
@@
  • Abnormal elastic scattering cross sections. Elastic scattering is affected by the coupling to non-elastic processes. In particular, when coupling to breakup channels is important, elastic scattering cross sections are depleted with respect to the case of tightly bound nuclei. Some other key signatures are the departure of the elastic cross section with respect to the Rutherford cross section at sub-Coulomb energies and the disappearance of the Fresnel peak at near-barrier energies in reactions induced by halo nuclei on heavy targets San08 ; Pie12 ; Cub12 ; Pes17 .
@@
One of the most successful models for describing reactions involving weakly-bound nuclei, which incorporates in a natural way the few-body structure of these systemas as well as its breakup is the Continuum Discretized Coupled-Channels (CDCC) Aus87 . The model and its recent extensions will be discussed in Sec. 3. Other models, tailored to specific processes will be discussed along this review.
It is the purpose of this paper to review some of the theories developed and applied for describing such nuclear reactions with weakly bound nuclei. As any review, the present one will be necessarily incomplete and possibly biased by the expertise and personal taste of the authors with the hope that the reader will obtain, at least, a flavour of how reaction theory has helped the interpretation of experiments with these nuclear species.
@@ 2 The impact of weak-binding on the elastic cross section: optical model approach
As in the case of well-bound nuclei, the natural framework to study elastic scattering is the Optical Model (OM), which consists in solving a two-body Schrödinger equation with an effective nucleus-nucleus potential, referred to as the optical model potential (OMP) Sat83 . The form of this OMP can be formally derived from the microscopic nucleus-nucleus interaction. In his seminal work, Feshbach Fesh58 showed that the OMP can be formally written as a sum of two terms, a bare potential, which is the expectation value of the nucleus-nucleus potential in the ground state of the projectile+target system, and an additional term, the polarization potential, which accounts for the effect of non-elastic channels (inelastic scattering, transfer, breakup, fusion, …) on the elastic scattering: Explicitly:
@@
The effect of weak binding of any of the colliding partners affects both the bare and polarization potentials. The ground-state of the projectile will exhibit an extended tail (as compared to the case of tightly-bound nuclei) and this will affect the form of $V_{00}$. In addition, the weak-binding will enhance certain couplings described by the $\mathbf{V}_{0}$ operator, very prominently the breakup channels, and this will also modify the polarization term. An example is the Coulomb polarization potential described in the Introduction.
@@
In the presence of strong absorption, elastic scattering angular distributions between composite nuclei display some common features regardless of the interacting nuclei. At near-barrier energies, elastic distributions present two distinct regions. At small angles, where the Coulomb interaction dominates, the cross section remains close to the Rutherford prediction, but oscillates around it. The amplitudes of the oscillations increases as the scattering angle increases but remain roughly 25% within the Rutherford prediction. This is the so-called “illuminated” region. After this maximum, the cross section drops rapidly, becoming much smaller than the Rutherford cross sections at large angles (the “shadow” region). An example of this behaviour can be seen in the elastic scattering data of $\alpha$ particles on lead, shown in the left panel of Fig. 2.
These features have been commonly interpreted by invoking classical models. The oscillatory behaviour in the illuminated region can be qualitatively interpreted within a Fresnel picture as an interference between distant pure Coulomb trajectories with closer trajectories affected by the nuclear attractive potential, while the shadow region can be understood as a consequence of the existence of a maximum angle in the classical deflection function. A complete understanding of the observed pattern requires also the introduction of refractive effects which arise in nuclear scattering as a result of absorption Fri89 .
All these effects are naturally accommodated within the OM. For example, the solid line in the left panel of Fig. 2 represents an OM fit of the data using a standard (complex) Woods-Saxon potential for the real and imaginary parts. Although the OM is not unique, one may find potentials with radii close to the sum of the geometrical radii of the interacting nuclei and diffuseness parameters close to that of the nuclear densities.
@@
This behaviour can be also described within the OM but with significant modifications with respect to the of standard parametrizations extracted from well-bound nuclei. In particular, one needs very diffuse potentials with long-range absorptive tails. The effect of this long-range absorption is twofold: (i) it damps or suppresses the elastic Coulomb amplitude at small angles (i.e., at very large impact parameters), and (ii) it reduces the nuclear amplitude. The result is that the forward cross-section not only becomes smaller overall, but also loses its Coulomb-nuclear oscillations.
In order to reproduce the elastic scattering data of weakly-bound nuclei within the OM framework using standard Woods-Saxon potentials, a very long-range imaginary part is needed. Table 1 lists the OMP parameters for 4He+208Pb and 6He+208Pb at the same incident energy (22 MeV). The radii of the real and imaginary parts are kept the same for both systems. To reproduce the data, the 6He OM requires a significantly larger imaginary diffuseness parameter. The result of the OM calculation for the 6He is compared with the data for this reaction in the right panel of Fig. 2.
@@
This property is a clear indication of the presence of reaction channels dominated by long-range couplings. Natural candidates are the breakup of the projectile and the neutron transfer (either one or two) from the projectile to the target nucleus. Both processes are of a peripheral nature, which explains the long-range imaginary potential. To identify the actual channels producing this effect, one should go beyond the single-channel OM scheme and incorporate those channels into the reaction framework. In the case of the breakup of the projectile, this can be efficiently accomplished with the Continuum-Discretized Coupled-Channels (CDCC) method, which is described in the next sections.
@@ 3.1 Brief resumè of the CDCC method
with $H_{\mathrm{proj}}=T_{\vec{r}}+V_{bx}$ the projectile internal Hamiltonian, $T_{\vec{r}}$ and $T_{\vec{R}}$ are kinetic energy operators, $V_{bx}$ the inter-cluster interaction and $U_{bA}$ and $U_{xA}$ are the core-target and valence-target optical potentials (complex in general) describing the elastic scattering of the corresponding $b+A$ and $x+A$ sub-systems, at the same energy per nucleon of the incident projectile. In the CDCC method the three-body wave function of the system is expanded in terms of the eigenstates of the Hamiltonian $H_{\mathrm{proj}}$ including both bound and unbound states. Since the latter form a continuum, a procedure of discretization is applied, consisting in representing this continuum by a finite and discrete set of square-integrable functions. In actual calculations, this continuum must be truncated in excitation energy and limited to a finite number of partial waves $\ell$ associated to the relative co-ordinate $\vec{r}$. Normalizable states representing the continuum should be obtained for each $\ell,j$ values. Two main methods are used for this purpose:
@@
$\phi_{n}(\vec{r})={u_{\ell_{n}j_{n}}(r)\over r}[Y_{\ell}(\hat{r})\otimes\chi_{s}]_{j_{n}m_{j_{n}}},$ (6)
where $n\equiv\{[k_{n},k_{n+1}]\ell_{n},s,j_{n},m_{j_{n}}\}$ specifies the $n$-th bin, with $[k_{n},k_{n+1}]$ the wavenumber interval of the bin, $\ell$ the valence-core orbital angular momentum, $s$ the valence spin, and $\vec{j}=\vec{\ell}+\vec{s}$ the total angular momentum. The symbol $\otimes$ denotes the angular momentum coupling. The radial part of the bin is obtained as a linear combination (i.e., a wave packet) of scattering states as
@@
The inclusion of the breakup channels in the expansion (8) will affect the elastic channel wavefunction through the coupling potentials $U_{0n}$ ($n>0$). The entrance flux, given by the norm of the plane wave associated with the entrance channel, will be distributed among the elastic and breakup channels and this will naturally reduce the elastic cross section with respect to the situation in which those breakup channels are omitted. This effect turns out to be essential for a correct description of the elastic cross sections, as illustrated in Fig.  3 for the d+58Ni reaction at 80 MeV. The figure depicts also the discretization escheme employed in these calculations, which comprises $\ell=0,2$ continuum states.
@@
where, for the internal states, $\phi_{n,J,M}$, we have included an additional subscript $J$ to write explicitly the projectile spin and hence $n$ is now an index to enumerate states with the same $J$. Also, in the former equation $\vec{R}$ is the relative coordinate between the projectile center of mass and the target (assumed by now to be structureless), see Fig. 4. The index $\beta$ gathers the quantum numbers compatible with a given total angular momentum $J_{T}$, $\beta\equiv\{L,J,n\}$, where $L$ (projectile-target orbital angular momentum) and $J$ both couple to the total spin of the three-body system $J_{T}$. The spin of the target is ignored for simplicity of the notation. Each of these sets is called channel. In particular, $\beta_{i}$ denote the channels compatible with the initial state of the system (typically, the ground state of the projectile and target nuclei).
@@
The radial coefficients, $\chi_{\beta,\beta_{i}}^{J_{T}}(R)$, from which the scattering observables are extracted, are calculated by inserting (12) in the Schrödinger equation, giving rise to a system of coupled differential equations:
$\left(-{\hbar^{2}\over 2\mu}{d^{2}\over dR^{2}}+{\hbar^{2}L(L+1)\over 2\mu R^{2}}+\varepsilon_{n}-E\right)\chi^{J_{T}}_{\beta,\beta_{i}}(R)$
@@
with the coupling potentials:
@@
where $K_{\beta}=\sqrt{2\mu(E-\epsilon_{n})}$, $S^{J_{T}}_{\beta,\beta_{i}}$ are the S-matrix elements, $T^{J_{T}}_{\beta,\beta_{i}}$ are the T-matrix elements, $H^{\pm)}(KR)$ are the ingoing (--) and outgoing ($+$) Coulomb functions and $F_{L}(KR)$ is the regular Coulomb function Sat83 . Because of the relation between the Coulomb functions, the two equations above are fully equivalent. In fact, the coefficients $S^{J_{T}}_{\beta,\beta_{i}}$ and $T^{J_{T}}_{\beta,\beta_{i}}$ are related by $S^{J_{T}}_{\beta,\beta_{i}}=\delta_{\beta,\beta_{i}}+2iT^{J_{T}}_{\beta,\beta_{i}}$. Scattering amplitudes and differential cross sections are expressed in terms of these coefficients (see, e.g., Glen04 ; Sat83 ).
Figure 5: Application of the CDCC method to 6Li+40Ca elastic scattering at 156 MeV. Experimental data are from Ref. Maj78 . Adapted from Mor09b .
The standard CDCC method is based on a strict three-body model of the reaction ($b+x+A$), and has proven rather successful to describe elastic and breakup cross sections of deuterons and other weakly bound two-body nuclei, such as 6,7Li and 11Be (see Fig. 5). However, it has limitations. The assumption of inert bodies is not always justified, since excitations of the projectile constituents ($b$ and $x$) and of the target ($A$) may take place along with the projectile dissociation. Furthermore, the two-body picture may be inadequate for some nuclei as, for example, in the case of the Borromean systems (e.g. 6He, 11Li). Some extensions of the CDCC method to deal with the these situations are outlined below.

### v2

@@ 1 Introduction
Nuclear reactions are key tools to extract information about the structure of atomic nuclei, and the dynamical phenomena arising from nucleus-nucleus forces. When one of the colliding partners is weakly bound, new phenomena and mechanisms arise, requiring in some cases specific reaction frameworks tailored to the peculiarities of these systems and their interactions.
Weakly-bound nuclei appear in the proximity of the neutron and proton driplines, a region where new exotic structures and phenomena are found. Promiment examples are halo nuclei, weakly bound nuclei composed of a compact core and one or two loosely bound nucleons with an unusually large matter radius, or Borromean systems, three-body systems with no bound binary subsystems, such as ${}^{9}\text{Be}~{}(\alpha+\alpha+n)$ or ${}^{6}\text{He}~{}(\alpha+n+n)$.
Although the field has experienced a great impulse in the last decades, the physics of nuclear scattering with weakly-bound projectiles is not new and many of the problems that are being addressed were already recognized much earlier as, for example, in the context of low-energy deuteron scattering. The deuteron, while being a stable nucleus, displays many halo-like features, such as weak binding ($E_{b}=2.22$ MeV), large spatial extension (the proton-neutron separation is about 3.8 fm) and no bound excited states.
To illustrate the role of the weak binding on the scattering observables, let us consider the scattering of low-energy deuterons (a few MeV) by a heavy target nucleus, like 208Pb. Initially, when the deuteron is far apart from the target nucleus, it is in an internal state given by the ground-state of the proton-neutron Hamiltonian. As the deuteron approaches the target, it will feel its Coulomb repulsion. If the deuteron were a point-like particle, the effect of this interaction would be to distort the trajectory of the deuteron, without altering its internal structure. However, this is not the case. The Coulomb interaction acts on the proton, whose distance from the center-of-mass of the deuteron is $\sim$2 fm. Consequently, in addition to the monopole Coulomb potential ($\propto 1/R$) the deuteron will feel higher-order Coulomb multipoles arising from the expansion, valid for $r\ll R$, given by Cle62b
@@
where $\vec{R}$ is the coordinate from the target to the deuteron c.m.  $\vec{r}$ the proton-neutron relative coordinate and $\theta$ is the angle between them (see Fig. 1). The main deviation from the point Coulomb interaction is caused by the dipole term which depends on the orientation of the proton-neutron relative coordinate with respect to the deuteron-target coordinate. Specifically, when the proton is closer to the target nucleus ($\cos(\theta)<0$), the dipole potential will add a positive contribution, whereas when the proton is farther from the c.m. of the deuteron the contribution will be negative. Classically, we may think of this problem as an electric dipole moving in a slowly varying electric field, in which the tidal force exerted on the dipole will favour the configuration of Fig. 1. Quantum-mechanically, the perturbation induced by the dipole force will couple the deuteron ground state (positive parity) with negative-parity states. Since the ground-state is the only bound state, these states necessarily appear in the continuum. Therefore, the polarization effect will modify the state of the system, producing a new one in which the ground state is mixed with continuum states of negative parity. This has a twofold effect on the outcome of the scattering process. First, it will produce deviations of the elastic scattering with respect to the Rutherford formula. Second, the coupling with the positive-energy states will give rise to some dissociation probability, that is, the breakup of the deuteron into a proton and a neutron. In the next sections, we will discuss several methods to incorporate these two effects in the reaction formalisms.
@@
  • Abnormal elastic scattering cross sections. Elastic scattering is affected by the coupling to non-elastic processes. In particular, when coupling to breakup channels is important, elastic scattering cross sections are depleted with respect to the case of tightly bound nuclei. Some other key signatures are the departure of the elastic cross section from the Rutherford cross section at sub-Coulomb energies and the disappearance of the Fresnel peak at near-barrier energies in reactions induced by halo nuclei on heavy targets San08 ; Pie12 ; Cub12 ; Pes17 .
@@
One of the most successful models for describing reactions involving weakly-bound nuclei, which incorporates in a natural way the few-body structure of these systems as well as its breakup, is the Continuum Discretized Coupled-Channels (CDCC) Aus87 . The model and its recent extensions will be discussed in Sec. 3. Other models, tailored to specific processes, will be discussed along this review.
It is the purpose of this paper to review some of the theories developed and applied for describing nuclear reactions with weakly bound nuclei. As any review, the present one will be necessarily incomplete and possibly biased by the expertise and personal taste of the authors but with the hope that the reader will obtain, at least, a flavour of how reaction theory has helped the interpretation of experiments with these nuclear species.
@@ 2 The impact of weak-binding on the elastic cross section: optical model approach
As in the case of well-bound nuclei, the natural framework to study elastic scattering is the Optical Model (OM), which consists in solving a two-body Schrödinger equation with an effective nucleus-nucleus potential, referred to as the optical model potential (OMP) Sat83 . The form of this OMP can be formally derived from the microscopic nucleus-nucleus interaction. In his seminal work, Feshbach Fesh58 showed that the OMP can be formally written as a sum of two terms, a bare potential, which is the expectation value of the nucleus-nucleus potential in the ground state of the projectile+target system, and an additional term, the polarization potential, which accounts for the effect of non-elastic channels (inelastic scattering, transfer, breakup, fusion, …) on elastic scattering: Explicitly:
@@
The effect of weak binding of any of the colliding partners affects both the bare and polarization potentials. The ground-state of the projectile will exhibit an extended tail (as compared to the case of tightly-bound nuclei) and this will influence the form of $V_{00}$. In addition, the weak-binding will enhance certain couplings described by the $\mathbf{V}_{0}$ operator, very prominently the breakup channels, and this will also modify the polarization term. In general, the evaluation of the Feshbach operator associated with the polarization potential is very involved and very often it is replaced by some phenomenological form. A few cases exist however in which a explicit evaluation is feasible, at least for specific nonelastic processes. One such example is the so-called adiabatic Coulomb polarization potential, briefly introduced in the next subsection.
2.1 Evaluation of the polarization potential in a simple case: the adiabatic polarization potential
In the introductory section, we described the case of deuteron scattering by a heavy-target nucleus. It was argued that the tidal force caused by the action of the Coulomb field on the proton produced a modification of the Coulomb-target interaction with respect to the point-Coulomb case [c.f. Eq. (1)]. Quantum-mechanically, the effect of this modified Coulomb interaction will be to induce coupling to deuteron breakup states. This is in fact one of the contributions included in the Feshbach polarization potential given by the second term of Eq. (3). This can be analytically evaluated in the so-called adiabatic limit, which assumes that the excitation energies are high enough so the characteristic time for a transition to a state is small compared to the characteristic time for the collision Cle62b . Applying second-order perturbation theory, one gets the following expression for this adiabatic polarization potential Cle62b ; Ald75 :
$V_{\mathrm{pol}}(R)=-\sum_{n\neq 0}\frac{|\langle n|V_{\mathrm{dip}}|0\rangle|^{2}}{E_{n}-E_{0}}=-\frac{1}{2}\alpha\frac{(Z_{T}e)^{2}}{R^{4}},$ (4)
It is to be noted that the adiabiatic polarization potential given by Eq. (4) is purely real and, as such, does not describe the effect of the deuteron breakup that would remove flux from the elastic channel. This point will be revisited in Sec. 7.2, where another version of the polarization potential, which accounts for deuteron breakup, will be presented.
The polarization potential (4) adds an attractive contribution to the Coulomb point-particle potential that, when used in the Schrödinder equation, will produce a small, but measurable deviation of the elastic cross section with respect to the well-known Rutherford formula. An accurate measurement of sub-Coulomb elastic scattering data can therefore be used to extract the dipole polarizability of the deuteron (and other polarizable systems). This idea was used by Rodning et al. Rod82 to infer the deuteron polarizability parameter from the analysis of sub-Coulomb deuteron elastic scattering on lead.
To avoid the determination of absolute cross sections, in their analysis the authors of Ref. Rod82 introduced the adimensional quantity
$R(E)=\frac{\sigma(E=3\text{ MeV},\theta_{1}=60^{\circ})}{\sigma(E=3\text{ MeV},\theta_{2}=150^{\circ})}\frac{\sigma(E,\theta_{2}=150^{\circ})}{\sigma(E,\theta_{1}=60^{\circ})}.$ (5)
which is identically unity for pure Rutherford scattering. The deuteron polarizability will give rise to values of $R(E)$ which are less than unity and the deviation increases with increasing scattering energy. In Fig. 2 we show the measured values of $R(E)$ from Rod82 along with two calculations. One in which only a nuclear potential is included, which plays the role of a bare part of Eq. (2) and a second calculation including the effect of the deuteron poloarizability in the adiabatic approximation, Eq. (4), with $\alpha=0.70\pm 0.05$ fm3, the value extracted in Rod82 by comparison with the data.
This constitutes a neat and beautiful example in which an important structure quantity can be inferred from reaction observables.
Figure 2: Experimental Rod82 values for the quantity $R(E)$, defined in Eq. (5) for deuteron scattering on 208Pb at sub-Coulomb energies. The dashed line is a single-channel calculation including only the nuclear potential. The solid line is the single-channel calculation include also the effect of the Coulomb dipole polarizability by means of the adiabatic polarization potential of Eq. (4). See text for details.
2.2 The phenomenological optical model
@@
In the presence of strong absorption, elastic scattering angular distributions between composite nuclei display some common features regardless of the interacting nuclei. At near-barrier energies, elastic distributions present two distinct regions. At small angles, where the Coulomb interaction dominates, the cross section remains close to the Rutherford formula prediction, but oscillates around it. The amplitudes of the oscillations increase as the scattering angle increases but remain roughly 25% within the Rutherford prediction. This is the so-called “illuminated” region. After this maximum, the cross section drops rapidly, becoming much smaller than the Rutherford cross sections at large angles (the “shadow” region). An example of this behaviour can be seen in the elastic scattering data of $\alpha$ particles on lead, shown in the left panel of Fig. 3.
These features have been commonly interpreted by invoking classical models. The oscillatory behaviour in the illuminated region can be qualitatively interpreted within a Fresnel picture as an interference between distant pure Coulomb trajectories with closer trajectories affected by the nuclear attractive potential, while the shadow region can be understood as a consequence of the existence of a maximum angle in the classical deflection function Sat83 . A complete understanding of the observed pattern requires also the introduction of refractive effects which arise in nuclear scattering as a result of absorption Fri89 .
All these effects are naturally accommodated within the OM. For example, the solid line in the left panel of Fig. 3 represents an OM fit of the data using a standard (complex) Woods-Saxon potential. Although the OM is not unique, one may find potentials with radii close to the sum of the geometrical radii of the interacting nuclei and diffuseness parameters close to those of the nuclear densities.
@@
This behaviour can also be described within the OM but with potential parameters very different from those used in standard parametrizations extracted from well-bound nuclei. In particular, one needs very diffuse potentials with long-range absorptive tails. The effect of this long-range absorption is twofold: (i) it damps or suppresses the elastic Coulomb amplitude at small angles (i.e., at very large impact parameters), and (ii) it reduces the nuclear amplitude. The result is that the forward cross-section not only becomes smaller overall, but also loses its Coulomb-nuclear oscillations.
In order to reproduce the elastic scattering data of weakly-bound nuclei within the OM framework using standard Woods-Saxon potentials, a very long-range imaginary part is needed. Table 1 lists the OMP parameters for 4He+208Pb and 6He+208Pb at the same incident energy (22 MeV). The radii of the real and imaginary parts are kept the same for both systems. To reproduce the data, the 6He OM requires a significantly larger imaginary diffuseness parameter. The result of the OM calculation for 6He is compared with the data for this reaction in the right panel of Fig. 3.
@@
This property is a clear indication of the presence of reaction channels dominated by long-range couplings. Natural candidates are the breakup of the projectile and neutron transfer (either one or two) from the projectile to the target nucleus. Both processes are of a peripheral nature, which explains the long-range imaginary potential. To identify the actual channels producing this effect one must go beyond the single-channel OM scheme and incorporate those channels into the reaction framework. In the case of the breakup of the projectile, this can be efficiently accomplished with the Continuum-Discretized Coupled-Channels (CDCC) method, which is described in the next sections.
@@ 3.1 Brief resumè of the CDCC method
with $H_{\mathrm{proj}}=T_{\vec{r}}+V_{bx}$ the projectile internal Hamiltonian, $T_{\vec{r}}$ and $T_{\vec{R}}$ are kinetic energy operators, $V_{bx}$ the inter-cluster interaction and $U_{bA}$ and $U_{xA}$ are the core-target and valence-target optical potentials (complex in general) describing the elastic scattering of the corresponding $b+A$ and $x+A$ sub-systems, at the same energy per nucleon of the incident projectile. In the CDCC method the three-body wave function of the system is expanded in terms of the eigenstates of the Hamiltonian $H_{\mathrm{proj}}$ including both bound and unbound states. Since the latter form a continuum, a procedure of discretization is applied, consisting in representing this continuum by a finite and discrete set of square-integrable functions. In actual calculations, this continuum must be truncated in excitation energy and limited to a finite number of partial waves associated with the relative co-ordinate $\vec{r}$. Normalizable states representing the continuum should be obtained for each set of orbital and total angular momenta. Two main methods are used for this purpose:
@@
$\phi_{n}(\vec{r})=\frac{u_{\ell_{n}j_{n}}(r)}{r}[Y_{\ell}(\hat{r})\otimes\chi_{s}]_{j_{n}m_{j_{n}}},$ (7)
where $n\equiv\{[k_{n},k_{n+1}]\ell_{n},s,j_{n},m_{j_{n}}\}$ specifies the $n$-th bin, with $[k_{n},k_{n+1}]$ the wavenumber interval of the bin, $\ell$ the valence-core orbital angular momentum, $s$ the valence spin, and $\vec{j}=\vec{\ell}+\vec{s}$ the total angular momentum. The symbol $\otimes$ denotes angular momentum coupling. The radial part of the bin is obtained as a linear combination (i.e., a wave packet) of scattering states as
@@
The inclusion of the breakup channels in expansion (9) will affect the elastic channel wavefunction through the coupling potentials $U_{0n}$ ($n>0$). The entrance flux, given by the norm of the plane wave associated with the entrance channel, will be distributed among the elastic and breakup channels and this will naturally reduce the elastic cross section with respect to the situation in which those breakup channels are omitted. This effect turns out to be essential for a correct description of the elastic cross sections, as illustrated in Fig. 4 for the d+58Ni reaction at 80 MeV. The figure also depicts the discretization scheme employed in these calculations, which comprises $\ell=0,2$ continuum states.
@@
where, for the internal states, $\phi_{n,J,M}$, we have included an additional subscript $J$ to write explicitly the projectile spin and hence $n$ is now an index to enumerate states with the same $J$. Also, in the former equation $\vec{R}$ is the relative coordinate between the projectile center of mass and the target (assumed by now to be structureless), see Fig. 5. The index $\beta$ gathers the quantum numbers compatible with a given total angular momentum $J_{T}$, $\beta\equiv\{L,J,n\}$, where $L$ (projectile-target orbital angular momentum) and $J$ both couple to the total spin of the three-body system $J_{T}$. The spin of the target is ignored for simplicity of notation. Each of these sets is called a channel. In particular, $\beta_{i}$ denotes the channels compatible with the initial state of the system (typically, the ground state of the projectile and target nuclei).
@@
The radial coefficients, $\chi_{\beta,\beta_{i}}^{J_{T}}(R)$, from which the scattering observables are extracted, are calculated by inserting Eq. (13) in the Schrödinger equation, giving rise to a system of coupled differential equations:
$\left(-\frac{\hbar^{2}}{2\mu}{d^{2}\over dR^{2}}+\frac{\hbar^{2}L(L+1)}{2\mu R^{2}}+\varepsilon_{n}-E\right)\chi^{J_{T}}_{\beta,\beta_{i}}(R)$
@@
where $\varepsilon_{n}$ is the nominal energy of function $\phi_{n}$ and with the coupling potentials:
@@
where $K_{\beta}=\sqrt{2\mu(E-\epsilon_{n})}$, $S^{J_{T}}_{\beta,\beta_{i}}$ are the S-matrix elements, $T^{J_{T}}_{\beta,\beta_{i}}$ are the T-matrix elements, $H^{\pm)}_{L}(KR)$ are the ingoing (--) and outgoing ($+$) Coulomb functions and $F_{L}(KR)$ is the regular Coulomb function Sat83 . Because of the relation between the Coulomb functions, the two equations above are fully equivalent. In fact, the coefficients $S^{J_{T}}_{\beta,\beta_{i}}$ and $T^{J_{T}}_{\beta,\beta_{i}}$ are related by $S^{J_{T}}_{\beta,\beta_{i}}=\delta_{\beta,\beta_{i}}+2iT^{J_{T}}_{\beta,\beta_{i}}$. Scattering amplitudes and differential cross sections are expressed in terms of these coefficients (see, e.g., Glen04 ; Sat83 ).
Figure 6: Application of the CDCC method to 6Li+40Ca elastic scattering at 156 MeV. The solid and dashed lines are the CDCC calculations with and without inclusion of the 6Li ($\alpha$+d) continuum. Experimental data are from Ref. Maj78 . Adapted from Mor09b
The standard CDCC method is based on a strict three-body reaction model ($b+x+A$), and has proven rather successful in describing elastic and breakup cross sections of deuterons and other weakly bound two-body nuclei, such as 6,7Li and 11Be (see Fig. 6). However, it has limitations. The assumption of inert bodies is not always justified, since excitations of the projectile constituents ($b$ and $x$) and of the target ($A$) may take place along with projectile dissociation. Furthermore, the two-body picture may be inadequate for some nuclei as, for example, in the case of the Borromean systems (e.g. 6He, 11Li). Some extensions of the CDCC method to deal with these situations are outlined below.